OpenStudy (anonymous):

Find the exact value of the expression. Do not use calculator. tan 60 degrees - cos(30) degrees/cos(60) degrees

OpenStudy (anonymous):



OpenStudy (yttrium):

\[\tan60 - \frac{ \cos30 }{ \cos60 } = \sqrt{3} - \frac{ \frac{ \sqrt{3} }{ 2 } }{ \frac{ 1 }{ 2 } } = \sqrt{3} - \sqrt{3} = 0\]

OpenStudy (anonymous):

oh ok thanks

OpenStudy (yttrium):

medal?